# HIGHT: Hierarchical Graph Tokenization for Graph-Language Alignment

### v1

@@ 5.2 Motif Hallucination
We first evaluate motif hallucination results of the LGLMs with node-centric and with HIGHT tokenization with MotifHallu. All the evaluated models only undergo the stage 1 instruction tuning to ensure a fair comparison. We have not included the other generalist baselines as we find they consistently answer “Yes”. In addition, in order to avoid the drawbacks that LGLMs may output answers that do not follow the instructions, we compare the loss values by feeding the answers of “Yes” and “No”, and take the one with a lower autoregressive language modeling loss as the answer. Following the practice in LVLMs, we present the F1 scores, accuracies, and the ratio that the model answers “Yes” (Li et al., 2023c). Given the severe imbalance of positive and negative samples in natural molecules, we separately report the F1 scores for positive and negative classes.
@@
The results are given in Table 1, which show that the LGLMs with node-centric tokenization consistently answer with “Yes” despite the absence of the corresponding functional groups. In contrast, HIGHT significantly improves the worst class hallucinations up to $40$% in terms of F1 scores, and the overall accuracies up to $30$%, thereby reducing the hallucination of LGLMs to the functional groups that do not exist in the molecule.
Method BACE $\uparrow$ BBBP $\uparrow$ HIV $\uparrow$
# Molecules 1513 2039 41127
Specialist Models
KV-PLM (Zeng et al., 2022) 78.5 70.5 71.8
GraphCL (You et al., 2020) 75.3 69.7 78.5
GraphMVP-C (Liu et al., 2022) 81.2 72.4 77.0
MoMu (Su et al., 2022) 76.7 70.5 75.9
MolFM (Luo et al., 2023a) 83.9 72.9 78.8
Uni-Mol (Zhou et al., 2023) 85.7 72.9 80.8
Galactica-6.7B (Taylor et al., 2022) 58.4 53.5 72.2
Galactica-30B (Taylor et al., 2022) 72.7 59.6 75.9
Galactica-120B (Taylor et al., 2022) 61.7 66.1 74.5
LLM Based Generalist Models
Vicuna-v1.5-13b-16k (4-shot) (Chiang et al., 2023) 49.2 52.7 50.5
InstructMol-G 63.2 55.4 57.5
InstructMol-GS 57.9 55.3 50.2
HIGHT-G 69.0 56.8 58.1
HIGHT-GS 68.4 59.4 58.6
Table 2: ROC-AUC Results of molecular property prediction tasks (classification) on MoleculeNet (Wu et al., 2017).
@@
We also conduct simple ablation studies by additionally incorporating the 1D sequence inputs with SELFIES following the literature (Fang et al., 2024; Cao et al., 2023). Contrary to previous results that additionally feeding the 1D sequence always improves the performance of LGLMs, We find that the additional 1D sequence may increase the degree of the hallucination. We suspect that it could be caused by the extremely long sequences of the SELFIES (Krenn et al., 2019) that may distract the attention signals of LLMs. Nevertheless, HIGHT suffers less from the distraction and maintains a better performance.
In addition, we also evaluate HIGHT without the tuning of HiPubChem. Aligned with our discussion in Sec. 3.2, HIGHT without HiPubChem will still suffer the hallucination, due to the low quality of the instruction tuning data. Interestingly, simply incorporating the hierarchical information at the architecture level can already help with the perception of graph information in LGLMs, which improves the robustness against hallucination, aligned with our discussion as in Sec. 4.1.
5.3 Molecular Property Prediction
In molecular property prediction, we leverage $3$ datasets BACE, BBBP, and HIV from MoleculeNet to evaluate the classification performance with ROC-AUC. We also adopt the regression based property prediction dataset from (Fang et al., 2024), where we evaluate several quantum chemistry measures such as HUMO, LUMO, and HUMO-LUMO gap (Ramakrishnan et al., 2014). The evaluation metric used to evaluate the regression based molecular property prediction is Mean Absolute Error (MAE). All the datasets are converted into instruction formats following previous works (Fang et al., 2024; Cao et al., 2023).
@@ 5.3 Molecular Property Prediction
HIGHT-GS 0.0056 0.0065 0.0077 0.0066
The results of molecular property prediction are given in Table 2 and Table 3 for classification and regression, respectively. We can find that, no matter for classification or regression-based molecular property prediction, HIGHT always significantly boosts the performance. The improvements brought by HIGHT serve as strong evidence verifying our discussion in Sec. 4 about the importance of hierarchical information for graph-language alignment. Nevertheless, there remains a gap between the specialist models in classification based molecular property prediction, leaving space for future improvements in terms of the adapters or instruction tuning configurations (Liu et al., 2023a).
@@
MolXPT (Liu et al., 2023f) 0.594 0.505 0.660 0.511 0.597 0.626
GIT-Mol-graph (Liu et al., 2024b) 0.290 0.210 0.540 0.445 0.512 0.491
GIT-Mol-SMILES (Liu et al., 2024b) 0.264 0.176 0.477 0.374 0.451 0.430
GIT-Mol-(graph+SMILES) (Liu et al., 2024b) 0.352 0.263 0.575 0.485 0.560 0.430
@@
HIGHT-GS 0.498 0.397 0.582 0.414 0.518 0.525
@@
HIGHT-GS 0.202 0.914 20.194 0.772 0.623 0.577 0.999
Table 5: Results of chemical reaction tasks. These tasks encompass reagent prediction, forward reaction prediction, and retrosynthesis. $\dagger$: few-shot ICL results from Fang et al. (2024). *: use task-specific instruction data to finetune.
5.4 Molecular Description Generation
For the task of molecular description generation or molecular captioning, we adopt the widely used benchmark ChEBI-20 (Edwards et al., 2021). Given the molecules, ChEBI-20 evaluates the linguistic distances of the generated molecule captions of molecular characteristics such as structure, properties, biological activities etc.. Following the common practice, we report the metrics of BLEU (Papineni et al., 2002), ROUGE (Lin, 2004) and Meteor (Banerjee and Lavie, 2005). The LGLMs are trained using the ChEBI-20 train split and evaluated using the test split. The final is selected according to the best training loss.
The results are given in Table 4. We can find that HIGHT consistently brings significant improvements over LGLMs with node-centric tokenization, again verifying our discussion about the importance of hierarchical graph information for graph-language alignment. Nevertheless, compared to the specialist models such as MoT5 (Edwards et al., 2022) that are pretrained on a significant amount of molecule-text related corpus, there remains a gap for generalist LGLMs even with HIGHT. The gap calls for interesting future investigations on how to properly incorporate HIGHT into the pretraining of the LGLMs.
5.5 Chemical Reaction Prediction
@@
For chemical reaction prediction tasks, we incorporate three tasks from Mol-Instructions (Fang et al., 2024), i.e., reagent prediction, forward reaction prediction, and retrosynthesis prediction, which are crucial for AI-aided drug discovery. Reagent prediction aims to predict the suitable reagents for a particular chemical reaction. Forward reaction prediction aims to predict the products of a chemical reaction, given the reactants and the reagents. Retrosynthesis prediction aims to predict the suitable reactants given a target product. The inputs and outputs for chemical reaction-related tasks adopt the SELFIES (Krenn et al., 2019) as recommended by (Fang et al., 2024). In terms of the evaluation metrics, we adopt both linguistic distance metrics such as BLEU (Papineni et al., 2002) and Levenshtein (Yujian and Bo, 2007), and molecular similarity measures such as similarity of the molecular fingerprints by RDKit (Landrum, 2016).
@@
The results are given in Table 5. It can be found that across all tasks in chemical reaction prediction, LGLMs with HIGHT consistently and significantly improve the performances compared to the node-centric tokenization. Meanwhile, LGLMs with HIGHT achieve the state-of-the-art results in several tasks and metrics, compared to other generalist models that even incorporate a stronger LLM backbone such as Mol-Instruction which adopts Llama-2 (Touvron et al., 2023b), and additional information of SELFIES.
@@ 6 Conclusions
This paper presents HIGHT, a novel hierarchical graph tokenization technique, which enhances the synergy between graph data, especially in molecular structures, and language models. By incorporating multi-level graph information, HIGHT improves the graph-language alignment performance, reducing hallucinations and boosting accuracy in molecular tasks, as validated through comprehensive benchmarking. Nevertheless, the current concentration on molecular graphs requires further verification for wider applicability to other forms of graph data, such as that originated from social networks. Despite the limitation, HIGHT represents a significant step forward in advancing language models’ graph comprehension capability, opening avenues for more refined cross-domain interactions and highlighting paths for future research in this direction.
@@
Appendix A Broader Impacts
This paper mainly focuses on how to best represent graph information for LLMs to understand better about the graphs. We demonstrate the effectiveness of our method on molecule-centric tasks which could facilitate the broader use of LLMs for tasks like AI-aided drug discovery and human-machine interactions in biomedicine. Besides, this paper does not raise any ethical concerns. This study does not involve any human subjects, practices to data set releases, potentially harmful insights, methodologies and applications, potential conflicts of interest and sponsorship, discrimination/bias/fairness concerns, privacy and security issues, legal compliance, and research integrity issues.
Appendix B Details of Instruction Tuning Datasets
@@
ChEBI-20 26,407 3,300 Molecules and the associated Chemical Entities of Biological Interest (ChEBI) [Hastings et al., 2015] annotations.
@@
B.1 Details of the PubChem Dataset
PubChem [^2] https://pubchem.ncbi.nlm.nih.gov is one of the largest public molecule database [Kim et al., 2022], and has been widely adopted by the alignment training of LGLMs [Liu et al., 2023d, c, Cao et al., 2023]. Our construction of PubChem predominantly follows Liu et al. [2023c]. We will briefly describe the main steps and interested readers may refer the details to [Liu et al., 2023c]:

### v2

@@ 5.2 Motif Hallucination
We begin with a proof-of-concept study with motif hallucination. We mainly compare LGLMs with node-centric to that with HIGHT tokenization with MotifHallu after stage 1 instruction tuning. For non-regression-based models, we include two state-of-the-art LGLMs GIMLET (Zhao et al., 2023) and Galactica (Taylor et al., 2022). We do not include the other regression-based models as we found they consistently answered “Yes”, making a nuanced F1 comparison less informative for them. To avoid the issue of format following, we compare the loss values by feeding the answers of “Yes” and “No” to the corresponding LLM, calculating the language modeling losses, and taking the one from “Yes” and “No” with a lower loss as the answer.
Reduction of hallucination. Due to the class imbalance issue in MotifHallu, we first report comprehensive metrics in Table 1. It can be found that HIGHT maintains great balance for both positive and negative classes compared to baselines. Especially, in terms of macro F1 scores that are averaged across classes, respectively, HIGHT demonstrates significant improvements up to $14$%.
@@
The results of the tokenization-focused comparison are given in Table 2. Following the practice in LVLMs, we present the F1 scores, accuracies, and the ratio that the model answers “Yes” (Li et al., 2023c). Given the imbalance of positive and negative samples, we separately report the F1 scores for different classes. It can be found that the LGLMs with node-centric tokenization consistently answer with “Yes” despite the absence of the corresponding functional groups. In contrast, HIGHT significantly reduces the worst class hallucination up to $40$% in terms of F1 scores, and improves the accuracies up to $30$%. The improvements are consistent and significant with both vicunna and llama2 LLM backbones.
@@
Ablations with different inputs and LLM backbones. We also conduct simple ablation studies by additionally incorporating the 1D sequence inputs with SELFIES (Fang et al., 2024; Cao et al., 2023). Contrary to previous results that additionally feeding the 1D sequence always improves the performance of LGLMs, we find that the additional 1D sequence may increase the degree of the hallucination. We suspect that it could be caused by the extremely long sequences of the SELFIES (Krenn et al., 2019) that may distract the attention signals of LLMs. Nevertheless, HIGHT still suffers less from the distraction and performs better.
In addition, when without HiPubChem (or with the HIGHT architecture), LGLMs will still suffer the hallucination, due to the low quality of the instruction tuning data, demonstrating the necessity of both components of HIGHT.
@@
5.3 Molecular-Centric Benchmarks
Molecular property prediction requires LGLMs to answer about particular properties given the molecule. We use $8$ datasets BACE, BBBP, HIV, SIDER, ClinTox, MUV, and Tox21 from MoleculeNet, and CYP450 from GIMLET (Zhao et al., 2023) to evaluate the classification performance with ROC-AUC. We also adopt the regression-based property prediction datasets from (Fang et al., 2024), where we evaluate several quantum chemistry measures such as HUMO, LUMO, and HUMO-LUMO gap (Ramakrishnan et al., 2014) via Mean Absolute Error (MAE).
The results of molecular property prediction are given in Table 3 and Table 4 for regression and classification, respectively. We can find that HIGHT always significantly boosts the performance in both types of tasks. Remarkably, in CYP450 (Zhao et al., 2023), HIGHT significantly outperforms the state-of-the-art model, demonstrating the advances of LGLM with hierarchical graph tokenization. Interestingly, Llama-2 (Touvron et al., 2023b) can match the state-of-the-art performance in HIV in a few-shot setting, while performing significantly worse in other datasets, for which we suspect some data contamination might exist.
Table 4: ROC-AUC Results of molecular property prediction tasks (classification) on MoleculeNet (Wu et al., 2017). Evaluation on InstructMol and HIGHT adopt the likelihood of the tokens of “Yes” and “No”. Most of the instruction tuning datasets are from GIMLET (Zhao et al., 2023). SIDER and ClinTox are converted following the MoleculeNet task description.
Method BACE $\uparrow$ BBBP $\uparrow$ HIV $\uparrow$ SIDER $\uparrow$ ClinTox MUV $\uparrow$ Tox21 $\uparrow$ CYP450 $\uparrow$
# Molecules 1,513 2,039 41,127 1,427 1,478 93,087 7,831 16,896
# Tasks 1 1 1 27 2 17 12 5
KV-PLM (Zeng et al., 2022) 78.5 70.5 71.8 59.8 84.3 61.7 49.2 59.2
GraphCL (You et al., 2020) 75.3 69.7 78.5 60.5 76.0 69.8 73.9 -
GraphMVP-C (Liu et al., 2022) 81.2 72.4 77.0 60.6 84.5 74.4 77.1 -
MoleculeSTM-G (Liu et al., 2023b) 80.8 70.0 76.9 61.0 92.5 73.4 76.9 -
MoMu (Su et al., 2022) 76.7 70.5 75.9 60.5 79.9 60.5 57.8 58.0
MolFM (Luo et al., 2023a) 83.9 72.9 78.8 64.2 79.7 76.0 77.2 -
Uni-Mol (Zhou et al., 2023) 85.7 72.9 80.8 65.9 91.9 82.1 78.1 -
Galactica-1.3B (Taylor et al., 2022) 57.6 60.4 72.4 54.0 58.9 57.2 60.6 46.9
Galactica-6.7B (Taylor et al., 2022) 58.4 53.5 72.2 55.9 78.4 - 63.9 -
Galactica-30B (Taylor et al., 2022) 72.7 59.6 75.9 61.3 82.2 - 68.5 -
Galactica-120B (Taylor et al., 2022) 61.7 66.1 74.5 63.2 82.6 - 68.9 -
GIMLET (Zhao et al., 2023) 69.6 59.4 66.2 - - 64.4 61.2 71.3
LLama-2-7b-chat (4-shot) (Touvron et al., 2023b) 76.9 54.2 67.8 - - 46.9 62.0 57.6
LLama-2-13b-chat (4-shot) (Touvron et al., 2023b) 74.7 52.8 72.4 - - 47.9 57.5 55.6
InstructMol-G 64.3 48.7 50.2 51.0 50.0 50.0 59.0 59.1
HIGHT-G 77.1 61.8 63.3 58.8 55.3 51.1 67.4 80.5
@@ 5.3 Molecular-Centric Benchmarks
MolXPT (Liu et al., 2023e) 0.594 0.505 0.660 0.511 0.597 0.626
GIT-Mol-graph (Liu et al., 2024c) 0.290 0.210 0.540 0.445 0.512 0.491
GIT-Mol-SMILES (Liu et al., 2024c) 0.264 0.176 0.477 0.374 0.451 0.430
GIT-Mol-(graph+SMILES) (Liu et al., 2024c) 0.352 0.263 0.575 0.485 0.560 0.430
@@
Molecular description requires the LGLMs to generate a caption of the molecule. We adopt the widely used benchmark ChEBI-20 (Edwards et al., 2021) which evaluates the linguistic distances of the generated molecule captions of molecular characteristics such as structure, properties, biological activities etc.. We report the metrics of BLEU (Papineni et al., 2002), ROUGE (Lin, 2004) and Meteor (Banerjee & Lavie, 2005). The LGLMs are trained using the ChEBI-20 train split, selected according to the best training loss, and evaluated using the test split.
As shown in Table 5, HIGHT consistently brings significant improvements over LGLMs with node-centric tokenization. Nevertheless, compared to the molecular foundation models such as MoT5 (Edwards et al., 2022) pretrained on a significant amount of molecule-text related corpus, there remains a gap for regression-based LGLMs even with HIGHT. The gap calls for future investigations on how to incorporate HIGHT into the pretraining of the LGLMs properly.
Table 6: Results of chemical reaction tasks. These tasks encompass reagent prediction, forward reaction prediction, and retrosynthesis. $\dagger$: few-shot ICL results from (Fang et al., 2024). *: use task-specific instruction data to finetune.
@@
Chemical reaction prediction requires the LGLMs to predict the results of the chemical reaction analysis, which are crucial for AI-aided drug discovery (Fang et al., 2024). Reagent prediction aims to predict the suitable reagents for a particular chemical reaction. Forward reaction prediction aims to predict the products of a chemical reaction, given the reactants and the reagents. Retrosynthesis prediction aims to predict the suitable reactants given a target product. The inputs and outputs for chemical reaction related tasks adopt the SELFIES (Krenn et al., 2019) as recommended by (Fang et al., 2024). We report both linguistic distance metrics such as BLEU (Papineni et al., 2002) and Levenshtein (Yujian & Bo, 2007), and molecular similarity measures such as similarity of the molecular fingerprints (Landrum, 2016).
@@
As shown in Table 6, across all tasks in chemical reaction prediction, LGLMs with HIGHT consistently and significantly improve the performances compared to the node-centric tokenization. Meanwhile, LGLMs with HIGHT achieve state-of-the-art results in several tasks and metrics, compared to other regression-based LGLMs that even incorporate a stronger LLM backbone such as Mol-Instruction, and additional information of SELFIES.
(a) Different training settings
(b) Zero-shot transfer
(c) Ablation variants of HIGHT
Figure 3: Ablation studies.
5.4 Empirical Analysis
Generalist capabilities. We follow the previous practice in training and evaluating generalist models (Liu et al., 2023a) and consider the two settings: a) As shown in Fig. 3(a), we first train the model with all chemical reaction prediction data by $3$ epochs to elicit the format following and the knowledge adaption capabilities of the LGLMs after stage 1. The models are named with “-All”; b) As shown in Fig. 3(b), we train the model with retrosynthesis task data and evaluate the zero-shot transfer performance on forward reaction prediction. Under both settings, we can find that HIGHT boosts the generalist capabilities significantly.
Computation overhead. In Appendix E.1, we report the computation overhead of pretraining and inference as well as tunable parameters of HIGHT and InstructMol. It can be found that, although HIGHT requires longer training time and relatively higher tunable parameters, the absolute values are not high. Moreover, during inference, as LLM latency consumes most of the computation, HIGHT can even reduce the inference latency by generating more concise answers.
Ablation studies. To better understand the effectiveness of distinct components in HIGHT, we conduct ablation studies that train InstructMol (Cao et al., 2023) with the laplacian positional encodings or with HiPubChem, as given in Fig. 3(c). We can find that, merely incorporating positional encoding or hierarchical instruction tuning is not sufficient to achieve the same performance as HIGHT. On the contrary, without a proper architecture design as HIGHT, LGLMs with previous node-centric tokenization with HiPubChem will confuse LLMs and even lead to degenerated downstream task performances. In addition, we also compare LGLMs with llama2 backbone. As shown in Fig. 3(a), HIGHT still significantly boosts the performance. More ablation studies are provided in Appendix E.
@@ 6 Conclusions
This paper presents HIGHT, a novel hierarchical graph tokenization technique. By incorporating the hierarchical graph information, HIGHT improves the molecule-language alignment performance, reducing hallucinations and boosting accuracy in molecular tasks. Nevertheless, the current focus on molecular graphs requires further verification for wider applicability to other forms of graph data, such as those originating from social networks. Despite the limitation, HIGHT represents a significant step forward in advancing graph comprehension capability of LLMs, and highlighting paths for future research in this direction.
Meanwhile, incorporating 3D information into the graph-language alignment is also a promising future direction, especially for broader scientific tasks such as single-cell modeling and understanding. For example, built upon HIGHT, one could design a new 3D tokenizer to accommodate 3D properties of motifs, scale up 3D data to include amino acids in proteins and certain recurrent structures in RNA sequences, incorporate 3D positional encoding, and curate instruction tuning data with 3D descriptive captions.
Acknowledgments
We thank the reviewers for their valuable comments. JC was supported by RGC Young Collaborative Research Grant No. C2005-24Y.
Impact Statement
This paper mainly focuses on how to best represent graph information for LLMs to better understand the graphs. We demonstrate the effectiveness of our method on molecule-centric tasks, which could facilitate the broader use of LLMs for tasks like AI-aided drug discovery and human-machine interactions in biomedicine. Besides, this paper does not raise any ethical concerns. This study does not involve any human subjects, practices, to data set releases, potentially harmful insights, methodologies and applications, potential conflicts of interest and sponsorship, discrimination/bias/fairness concerns, privacy and security issues, legal compliance, and research integrity issues.
@@ Acknowledgments
Appendix A More Future Works
Built upon HIGHT, there are several promising future directions. For example, one could extend this study to more types of graphs, such as social networks and knowledge graphs, by exploring the crucial substructures therein:
  • Indeed, motifs generically exist in other types of graphs and are crucial for a variety of tasks (Ribeiro et al., 2021). For example, cliques can define boundaries between groups of people in social networks (Doreian & Woodard, 1994). The idea of HIGHT could be seamlessly applied to other graphs where we have some prior knowledge about critical motifs.
  • Meanwhile, when we do not have prior knowledge about the motifs, the GNNs intrinsically model the hierarchical nature of graphs in different orders (Ying et al., 2018) and thus can be integrated into LGLMs to learn the hierarchical graph information. A similar idea has been verified successful in graph transformers (Zhang et al., 2022).
  • Furthermore, one could also adopt interpretable GNNs to identify the critical subgraphs for the task (Yu et al., 2021; Miao et al., 2022; Chen et al., 2024b) that capture the underlying causal information about the underlying tasks (Chen et al., 2022, 2023; Yao et al., 2024; Liu et al., 2024a; Yao et al., 2025; Xu et al., 2025). It is also interesting to further investigate the hallucinations caused by the spurious correlations during the alignment (Wang et al., 2024).
Appendix B Comparison between other LGLMs
Table 7: Comparison between other LGLMs in terms of the backbone, instruction tuning, downstream usage for Molecular Property prediction, and capable tasks. It can be found that HIGHT is capable of various tasks, given limited pre-training data and information. Note that compared to the instruction tuning data for other LGLMs, such as KV-PLM (Zeng et al., 2022), which consists of papers with detailed information about molecules, the text descriptions in HIGHT contain relatively simple sentences.
Model Backbone Information Instruction Tuning Data Downstream # Tasks
HIGHT/InstructMol (Cao et al., 2023) GNN+Llama 2D Graph+Text HiPubChem-295k LoRA 7
GraphCL (You et al., 2020) GNN 2D Graph Downstream training data Finetuning 1
GraphMVP (Liu et al., 2022) GNN 2D+3D graph Geom-50K Finetuning 1
MoleculeSTM (Xia et al., 2023) GNN+SciBERT 2D Graph+3D Graph+Text PubChemSTM-280K Finetuning 6
KV-PLM (Zeng et al., 2022) BERT 1D SMILES+Text S2orc-300K academic papers Finetuning 5
MolT5 (Edwards et al., 2022) T5 1D SMILES+Text C4+ZINC(100M) Finetuning 2
Text+Chem T5 (Christofidellis et al., 2023) T5 1D SMILES+Text Multi task-33.5M Finetuning 5
MoMu (Su et al., 2022) GNN+BERT 2D Graph+Text Graph-Docuemnt Pair-15.6K Finetuning 4
MolFM (Luo et al., 2023a) GNN+BERT Knowledge Graph+2D+3D Graph+Text KG-15K+S2ORC-37M Finetuning 4
Uni-Mol (Zhou et al., 2023) Transformer 3D Graph molecule-209M+protein-3.2M Finetuning 4
Galactica (Taylor et al., 2022) GPT 1D SMILES/Text Documents-59M+chemicla prompts-2.5M zero-shot 12
MolXPT (Liu et al., 2023e) GPT2 1D SMILES+Text Mixed text-68M Finetuning 2
GIMLET (Zhao et al., 2023) GNN+T5 2D Graph+Text ChemBL-730K zero-shot 2
Appendix C Details of Instruction Tuning Datasets
@@
MoleculeNet-SIDER 1,141 143 Question answering about the ability of the side effects.
MoleculeNet-ClinTox 1,188 148 Question answering about the toxicology.
MoleculeNet-MUV 74,469 9,309 Question answering about PubChem bioAssay
MoleculeNet-Tox21 6,877 860 Question answering about Toxicology in the 21st century
CYP45- 13,516 1,690 Question answering about CYP PubChem BioAssay CYP 1A2, 2C9, 2C19, 2D6, 3A4 inhibition.
@@
ChEBI-20 26,407 3,300 Molecules and the associated Chemical Entities of Biological Interest (ChEBI) (Hastings et al., 2015) annotations.
@@
Property Prediction CYP450 https://www.nlm.nih.gov/web_policies.html The data is from Zhao et al. (2023) that curates PubChem BioAssay CYP 1A2, 2C9, 2C19, 2D6, 3A4 inhibition. Thus it shares the same license as PubChem.
@@
Table 10: Summary of inputs and outputs of the tasks in experiments.
input output
motif hallucination molecule and question about the existence of a motif yes or no
molecular property prediction (classification) molecule and question about the existence of the property yes or no
molecular property prediction (regression) molecule and question about the value of the property property value
molecular caption molecule and question asking for the molecular caption molecular caption
chemical reaction prediction molecules and question about the reaction molecular results
C.1 Details of the PubChem Dataset
PubChem [^3] https://pubchem.ncbi.nlm.nih.gov is one of the largest public molecule database (Kim et al., 2022), and has been widely adopted by the alignment training of LGLMs (Liu et al., 2023c, b; Cao et al., 2023). Our construction of PubChem predominantly follows Liu et al. (2023b). We will briefly describe the main steps and interested readers may refer the details to (Liu et al., 2023b):